# Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$

### v3

@@ 3. Classification of orbits
In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, which will complete a classification of coadjoint orbits for type $F_{4}$. For the next stage, we will need a concept of a singular root.
@@
It is obvious that the singular roots from $S_{\gamma}$ are divided into pairs of the form $\alpha+\beta=\gamma$.
Next, unless stated otherwise, everything will take place over the closure of the field $\overline{\mathbb{F}}$. We introduce two $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$-modules:
@@
Here $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ denotes the ring of polynomials over $\overline{\mathbb{F}}$ in independent variables in $x_{\alpha},\alpha\in\Phi^{+}$. Next, we define the pairing $M^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ by putting
@@
Everywhere below we will write $\lambda(x)$ instead of $\langle\lambda,x\rangle$ for $\lambda\in M^{*},x\in M$. Notice that $\mathfrak{n}$ is a $\mathbb{F}$-subspace in $M$, and $\mathfrak{n}^{*}$ is a $\mathbb{F}$-subspace in $M^{*}$. Therefore, two other pairings $\mathfrak{n}^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $M^{*}\times\mathfrak{n}\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ are automatically defined. We introduce a Lie algebra structure on $M$ in an obvious way: for $f,g\in\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $x,y\in\mathfrak{n}$, then
@@
Since $\mathfrak{n}$ can be identified with upper triangular matrices with coefficients from $\mathbb{F}$, we can identify $M$ with upper triangular matrices with coefficients from $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$. Therefore, we can define $\exp(x)$ for $x\in M$. Since the characteristic of $\mathbb{F}$ is sufficiently large, $\exp(x)$ is well defined for all $x\in M$. Let $N_{M}$ be the group consisting of the matrices $\{\exp(x),x\in M\}$. The actions $N_{M}\curvearrowright M$ and $N_{M}\curvearrowright M^{*}$ are naturally defined. The explicit description of the action $N_{M}\curvearrowright M^{*}$ is as follows: if $x\in M,\lambda\in M^{*}$, then
@@
Denote the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$. For forms $\lambda\in\mathfrak{n}^{*}$, we will consider the elements $\exp(\chi).\lambda\in M^{*}$. Let us designate $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$. Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ for certain $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
Now we formulate one of the most important theorems of this work. Let $y_{\beta},\beta\in\Phi^{+}$, be independent variables.
Theorem 3.2. Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, $\gamma\in\Phi^{+}$ satisfies condition (1). Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
@@
Example 3.3. Consider the same support as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that for $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds. Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}$. Let us explicitly write down the coordinates we are interested in. As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we will denote the structural constants in the expression $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. We have
@@
In fact, the validity of formula (2) in this example could have been understood without explicitly writing out the coordinates. It is enough to look at the Table 1. Due to the fact that $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, $x_{\alpha_{3}}$ can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Further, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$. Since $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ contains only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function in $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
Now let us move on to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. First, we will prove it in a particular case, for forms with a support consisting of one root. Obviously, any such a form lies in $S$.
@@
Proof. Denote $D=\textup{supp}(\lambda)=\{\delta\}$. Let us note that in this case, the equality (1) only depends on $D$ and not on $\lambda$, and holds for those $\gamma$ that are not singular to $\delta$. Let $\mu=\exp(\chi).\lambda$. We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}. Note that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$, so for these cases, the condition is obviously satisfied. It remains to prove it for $\gamma\prec\delta$.
@@
where $C\neq 0$ and $F(x)$ is a polynomial in $x_{\alpha}$, where $\alpha<\delta-\beta$. Thus,
@@
Let $\beta_{1},\beta_{2},\ldots,\beta_{2k}$ be the singular roots to $\delta$, listed in decreasing order with respect to lexicographical order. They are paired such that $\beta_{i}+\beta_{2k-i+1}=\delta$. Let $\gamma\notin S_{\delta}$ with $\delta\succ\gamma$. Consider $\mu(e_{\gamma})$. Suppose that there is no rational function $H$ over $\overline{\mathbb{F}}$ of independent variables $t_{\beta},\beta\succ\gamma$ such that $\mu(e_{\gamma})=H(\mu(e_{\beta}),\beta\succ\gamma)$.
@@
The coordinate $\mu(e_{\gamma})$ is a polynomial in $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$. It does not contain $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$. Let $\beta_{2k-j+1}$ be the largest root from $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ is in $\mu(e_{\gamma})$. We express $x_{\beta_{2k-j+1}}$ using formula (3) through $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}. Thus, now $\mu(e_{\gamma})$ is a polynomial that does not include $x_{\beta_{2k-j+1}}$. We perform a similar operation for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$. We note that at each step, $x_{\beta_{i}}$, which was initially absent and has already been removed, cannot reappear, as we are proceeding in decreasing order of singular roots. As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}(t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$. Note that all $\beta_{i}\succ\gamma,i=1,\ldots,j$.
@@
According to our assumption, $\mu(e_{\gamma})$ cannot be expressed in terms of the subsequent coordinates, which means that this polynomial is not identically equal to some number and contains at least one variable $x_{\alpha}$. Then, there exists an open Zariski subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is not a constant. Fix any such set $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$. Choose $\varepsilon\notin S_{\delta}$ and $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$ such that the polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is not constant. The field is algebraically closed, therefore for any $c_{0}\in\overline{\mathbb{F}}$, there exists such $c_{\varepsilon}\in\overline{\mathbb{F}}$ that $g(c_{\varepsilon})=c_{0}$. We fix $c_{0}$ and $c_{\varepsilon}$. We can choose any $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$. By substituting $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into the formula (3), we will sequentially assign values to $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ so that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$. For these values, $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
@@
Thus, there exists an open Zariski subset $V\subseteq\overline{\mathbb{F}}^{2k}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ there exist $\overline{\mu}\in\overline{N}.\lambda$ for which the coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values. The orbit $\overline{N}.\lambda$ is an algebraic variety. From our construction, it follows that there is a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, therefore $\dim\overline{N}.\lambda>2k$. But $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$. A variety of dimension $2k$ cannot contain $\overline{\mathbb{F}}^{2k+1}$. This is a contradiction. $\square$
We introduce more notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
@@
Now consider the following construction (cf. the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$). Let $\lambda\in S,~D=\textup{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1). Starting from $D$ and $\gamma$, we will construct the set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function in $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
Example 3.5. Table 1 helps to visualize how the polynomials $T_{\alpha,\beta}(x)$ look like. It is enough to look at the cell in the row numbered $\alpha$ and the column numbered $\beta$. For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, we denote the non-zero constant $C_{\gamma,\delta}$ by letting $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$. Each monomial in $T_{\alpha,\beta}(x)$ has the form
@@
where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 1 to $n$, $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$. Hence, it is enough to subsequently consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots. For instance, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
Note that if $\alpha\nleq\beta$ then $T_{\alpha,\beta}(x)=0$, while $T_{\alpha,\alpha}(x)=1$. In all other cases, $T_{\alpha,\beta}(x)$ has non-zero degree. Let us call a pair of roots $(\alpha,\beta)$ nontrivial if $\alpha<\beta$ (i.e., $T_{\alpha,\beta}(x)\neq const$). Obviously, each trivial pair belongs to all $\zeta^{\prime}_{D,\gamma}$.
Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$ again. There $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ do not contain nontrivial pairs. Indeed, there are no $\beta$ in $D$ such that $\beta>\alpha_{2}+2\alpha_{3}$ or $\beta>\alpha_{2}+\alpha_{3}+\alpha_{4}$. Since

### v4

@@ 3. Classification of orbits
In this section, we will prove Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, thereby completing the classification of coadjoint orbits for type $F_{4}$. For the next stage, we will need the concept of a singular root.
@@
It is obvious that the singular roots in $S_{\gamma}$ are divided into pairs of the form $\alpha+\beta=\gamma$.
From now on, unless stated otherwise, everything will take place over the algebraic closure $\overline{\mathbb{F}}$. We introduce two $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$-modules:
@@
Here $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ denotes the ring of polynomials over $\overline{\mathbb{F}}$ in the independent variables $x_{\alpha},\alpha\in\Phi^{+}$. We define a pairing $M^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ by
@@
In what follows, we will write $\lambda(x)$ instead of $\langle\lambda,x\rangle$ for $\lambda\in M^{*},x\in M$. Note that $\mathfrak{n}$ is a $\mathbb{F}$-subspace of $M$, and $\mathfrak{n}^{*}$ is a $\mathbb{F}$-subspace of $M^{*}$. Therefore, two other pairings, $\mathfrak{n}^{*}\times M\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $M^{*}\times\mathfrak{n}\to\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$, are automatically induced. We introduce a Lie algebra structure on $M$ in the obvious way: for $f,g\in\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$ and $x,y\in\mathfrak{n}$, set
@@
Since $\mathfrak{n}$ can be identified with upper-triangular matrices with entries in $\mathbb{F}$, we can identify $M$ with upper-triangular matrices whose entries lie in $\overline{\mathbb{F}}[x_{\alpha},\alpha\in\Phi^{+}]$. Consequently, we can define $\exp(x)$ for $x\in M$. Since the characteristic of $\mathbb{F}$ is sufficiently large, $\exp(x)$ is well-defined for all $x\in M$. Let $N_{M}$ be the group consisting of the matrices $\{\exp(x),x\in M\}$. The actions $N_{M}\curvearrowright M$ and $N_{M}\curvearrowright M^{*}$ are defined naturally. An explicit description of the action $N_{M}\curvearrowright M^{*}$ is as follows: for $x\in M$ and $\lambda\in M^{*}$,
@@
Denote by $\chi$ the element $\chi=\sum\limits_{\alpha\in\Phi^{+}}x_{\alpha}\otimes e_{\alpha}\in M$. For a form $\lambda\in\mathfrak{n}^{*}$, we will consider the element $\exp(\chi).\lambda\in M^{*}$. Let us also set $\overline{\mathfrak{n}}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}\mathfrak{n}$, $\overline{N}=\overline{\mathbb{F}}\otimes_{\mathbb{F}}N$. Note that if $\widetilde{x}=\sum\limits_{\alpha\in\Phi^{+}}\widetilde{x}_{\alpha}e_{\alpha}\in\overline{\mathfrak{n}}$ with $\widetilde{x}_{\alpha}\in\overline{\mathbb{F}}$, then $\exp(\widetilde{x}).\lambda=(\exp(\chi).\lambda)(\widetilde{x}_{\alpha},\alpha\in\Phi^{+})$.
Now we formulate one of the most important theorems of this work. Let $y_{\beta},\beta\in\Phi^{+}$ be independent variables.
Theorem 3.2. Let $\lambda\in S$, $D=\textup{supp}(\lambda)$, and let $\gamma\in\Phi^{+}$ satisfy condition (1). Then there exists a rational function $F_{D,\gamma}\in\overline{\mathbb{F}}(y_{\beta},\beta\succ\gamma)$ such that
@@
Example 3.3. Consider the same support as in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. Let $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. We will show that for this $\lambda$ and for all roots $\gamma$ satisfying condition (1), formula (2) holds. Let $\mu=\exp(\chi).\lambda,~\lambda=\sum\limits_{\alpha\in\Phi^{+}}\lambda_{\alpha}e_{\alpha}^{*},~\mu=\sum\limits_{\alpha\in\Phi^{+}}\mu_{\alpha}\otimes e_{\alpha}^{*}$. Let us write down explicitly the coordinates of interest. As in Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$, we denote the structural constants in the relation $Ce_{\alpha+\beta}=[e_{\alpha},e_{\beta}]$ by $C_{\alpha,\beta}$. Their values can be found in [KP23]. We have
@@
In fact, the validity of formula (2) in this example could have been understood without writing out the coordinates explicitly. It is enough to examine Table 1. Since $\lambda(e_{\alpha_{2}+2\alpha_{3}})=\mu(e_{\alpha_{2}+2\alpha_{3}})$ and $\mu(e_{\alpha_{2}+\alpha_{3}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, the variable $x_{\alpha_{3}}$ can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Consequently, since $\mu(e_{\alpha_{2}})-\lambda(e_{\alpha_{2}})$ depends only on $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{2}+2\alpha_{3}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{2}+2\alpha_{3}})$. Furthermore, $\mu(e_{\alpha_{3}+\alpha_{4}})=\lambda(e_{\alpha_{3}+\alpha_{4}})$. As $\mu(e_{\alpha_{4}})-\lambda(e_{\alpha_{4}})$ involves only $x_{\alpha_{3}}$ and $\lambda(e_{\alpha_{3}+\alpha_{4}})$, it can be expressed as a rational function of $\mu(e_{\alpha_{2}+\alpha_{3}})$ and $\mu(e_{\alpha_{3}+\alpha_{4}})$.
Now let us proceed to the proof of Theorem Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. First, we prove it in the special case of forms whose support consists of a single root. Any such form obviously belongs to $S$.
@@
Proof. Denote $D=\textup{supp}(\lambda)=\{\delta\}$. Note that in this case, condition (1) depends only on $D$, not on the specific $\lambda$, and it holds for those $\gamma$ that are not singular to $\delta$. Let $\mu=\exp(\chi).\lambda$. We need to prove that for any $\gamma\notin S_{\delta}$, the coordinate $\mu(e_{\gamma})$ is a function of the coordinates $\mu(e_{\beta}),\beta\succ\gamma$}. Observe that if $\gamma\succ\delta$, then $\mu(e_{\gamma})=0$, and if $\gamma=\delta$, then $\mu(e_{\delta})=\lambda(e_{\delta})$. For these two cases, the condition is obviously satisfied. It remains to prove it for $\gamma\prec\delta$.
@@
where $C\neq 0$ and $F(x)$ is a polynomial in variables $x_{\alpha}$ with $\alpha<\delta-\beta$. Consequently,
@@
Let $\beta_{1},\beta_{2},\ldots,\beta_{2k}$ be the singular roots to $\delta$, listed in decreasing order with respect to lexicographical order. They come in pairs such that $\beta_{i}+\beta_{2k-i+1}=\delta$. Now take $\gamma\notin S_{\delta}$ with $\delta\succ\gamma$ and consider $\mu(e_{\gamma})$. Suppose, for contradiction, that there is no rational function $H$ over $\overline{\mathbb{F}}$ in the independent variables $t_{\beta},\beta\succ\gamma$ such that $\mu(e_{\gamma})=H(\mu(e_{\beta}),\beta\succ\gamma)$.
@@
The coordinate $\mu(e_{\gamma})$ is a polynomial in the $x_{\beta}$ with coefficients from $\overline{\mathbb{F}}$. It does not contain the variable $x_{\beta_{2k-i+1}}$ if $\gamma\succ\beta_{i}$. Let $\beta_{2k-j+1}$ be the largest root in $S_{\delta}$ such that $x_{\beta_{2k-j+1}}$ appears in $\mu(e_{\gamma})$. Using formula (3), we express $x_{\beta_{2k-j+1}}$ in terms of $\mu(e_{\beta_{j}})$ and $x_{\alpha},\alpha<\beta_{2k-j+1}$}. Thus, $\mu(e_{\gamma})$ becomes a polynomial that no longer involves $x_{\beta_{2k-j+1}}$. We perform a similar elimination successively for $\beta_{2k-j+2},\beta_{2k-j+3}$, and so on, until $\beta_{2k}$. Note that at each step, a variable $x_{\beta_{i}}$ that was initially absent or already removed cannot reappear, because we proceed in decreasing order of the singular roots. As a result, $\mu(e_{\gamma})$ is represented as a polynomial that does not contain $x_{\beta},\beta\in S_{\delta}$, i.e., $\mu(e_{\gamma})=G(\mu(e_{\beta_{i}}),i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ for some polynomial $G\in\overline{\mathbb{F}}[t_{\beta_{i}},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta}]$. Observe that $\beta_{i}\succ\gamma$ for all $i=1,\ldots,j$.
@@
By our assumption, $\mu(e_{\gamma})$ cannot be expressed solely in terms of the subsequent coordinates. This implies that $G$ is not identically constant and actually depends on at least one variable $x_{\alpha}$. Then there exists a non-empty Zariski-open subset $U\subseteq\overline{\mathbb{F}}^{j}$ such that for any $c=(c_{1},c_{2},\ldots,c_{j})\in U$, the specialized polynomial $G(c_{i},i=1,\ldots,j;~x_{\alpha},\alpha\notin S_{\delta})$ is non-constant. Fix such a tuple $c_{1},c_{2},\ldots,c_{j}\in\overline{\mathbb{F}}$. Choose an index $\varepsilon\notin S_{\delta}$ and constants $c_{\alpha}\in\overline{\mathbb{F}},\alpha\notin S_{\delta},\alpha\neq\varepsilon$, so that the one-variable polynomial $g(x_{\varepsilon})=G(c_{i},i=1,\ldots,j;~x_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta},\alpha\neq\varepsilon)$ is non-constant. The field is algebraically closed, for any chosen $c_{0}\in\overline{\mathbb{F}}$ there exists a $c_{\varepsilon}\in\overline{\mathbb{F}}$ such that $g(c_{\varepsilon})=c_{0}$. Fix $c_{0}$ and the corresponding $c_{\varepsilon}$. Now, choose arbitrary values $c_{j+1},\ldots,c_{2k}\in\overline{\mathbb{F}}$. Substituting the values $c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta}$ into formula (3), we can determine values for $x_{\beta_{2k}},x_{\beta_{2k-1}},\ldots,x_{\beta_{1}}$ sequentially, in such a way that $\mu(e_{\beta_{1}})=c_{1},\mu(e_{\beta_{2}})=c_{2},\ldots,\mu(e_{\beta_{2k}})=c_{2k}$. For this assignment, we obtain $\mu(e_{\gamma})=G(c_{i},i=1,\ldots,j;~c_{\varepsilon},c_{\alpha},\alpha\notin S_{\delta})=c_{0}$.
@@
Therefore, there exists a non-empty Zariski-open subset $V\subseteq\overline{\mathbb{F}}^{2k+1}$ such that for any $(c_{0},c_{1},\ldots,c_{2k})\in V$ , we can find an element $\overline{\mu}\in\overline{N}.\lambda$ whose coordinates $\overline{\mu}(e_{\gamma}),\overline{\mu}(e_{\beta_{1}}),\ldots,\overline{\mu}(e_{\beta_{2k}})$ take these values. The orbit $\overline{N}.\lambda$ is an algebraic variety. Our construction yields a surjection from $\overline{N}.\lambda$ to $\overline{\mathbb{F}}^{2k+1}$, which implies $\dim\overline{N}.\lambda\geq 2k+1$. However, $\dim\overline{N}.\lambda=\dim\overline{\mathfrak{n}}.\lambda=\dim\mathfrak{n}.\lambda=|S_{\delta}|=2k$. This contradiction completes the proof. $\square$
We introduce further notation. Let $\alpha,\beta\in\Phi^{+}$. Denote
@@
Now consider the following construction (cf. the construction of the set $\zeta_{D,\gamma}$ used in the proof of the sufficient condition in Proposition Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$). Let $\lambda\in S,~D=\textup{supp}(\lambda)$, and let the root $\gamma\in\Phi^{+}$ satisfy condition (1). Starting from $D$ and $\gamma$, we construct a set $\zeta^{\prime}_{D,\gamma}\subset\Phi^{+}\times\Phi^{+}$: by definition, it consists of those pairs $(\alpha,\beta)$ for which the polynomial $T_{\alpha,\beta}(x)$ can be expressed as a rational function of $\{(\exp(\chi).\lambda)(e_{\delta})\}_{\delta\succ\gamma}$.
Example 3.5. Table 1 helps visualize the structure of the polynomials $T_{\alpha,\beta}(x)$. One simply examines the cell in row $\alpha$ and column $\beta$. For each pair $\gamma,\delta\in\Phi^{+}$ such that $\gamma+\delta\in\Phi^{+}$, denote the corresponding non-zero structure constant by $C_{\gamma,\delta}$, so that $e_{\gamma},e_{\delta}]=C_{\gamma,\delta}e_{\gamma+\delta}$. Each monomial in $T_{\alpha,\beta}(x)$ has the form
@@
where $\gamma_{1},\ldots,\gamma_{n}\in\Phi^{+},~\alpha+\gamma_{1}+\ldots+\gamma_{i}\in\Phi^{+}$ for all $i$ from 11 to $n$, and $\alpha+\gamma_{1}+\ldots+\gamma_{n}=\beta$. Hence, it suffices to consider all possible decompositions of $\beta-\alpha$ into a sum of positive roots. For example, $T_{\alpha_{3},\alpha_{2}+2\alpha_{3}+\alpha_{4}}(x)=-x_{\alpha_{2}+\alpha_{3}+\alpha_{4}}+\dfrac{1}{2}x_{\alpha_{2}}x_{\alpha_{3}+\alpha_{4}}+\dfrac{3}{2}x_{\alpha_{4}}x_{\alpha_{2}+\alpha_{3}}-\dfrac{2}{3}x_{\alpha_{2}}x_{\alpha_{3}}x_{\alpha_{4}}$.
Note that if $\alpha\nleq\beta$, then $T_{\alpha,\beta}(x)=0$, while $T_{\alpha,\alpha}(x)=1$. In all other cases, $T_{\alpha,\beta}(x)$ has positive degree. We call a pair $(\alpha,\beta)$ nontrivial if $\alpha<\beta$ (i.e., $T_{\alpha,\beta}(x)\neq const$). Obviously, each trivial pair belongs to all $\zeta^{\prime}_{D,\gamma}$.
Let us refer to Example Classification of coadjoint orbits for the maximal unipotent subgroup in the simple group of type $F_{4}$. There $D=\textup{supp}(\lambda)=\{\alpha_{4},\alpha_{3}+\alpha_{4},\alpha_{2},\alpha_{2}+2\alpha_{3}\}$. The sets $\zeta^{\prime}_{D,\alpha_{2}+2\alpha_{3}}$ and $\zeta^{\prime}_{D,\alpha_{2}+\alpha_{3}+\alpha_{4}}$ contain no nontrivial pairs. Indeed, there are no $\beta$ in $D$ such that $\beta>\alpha_{2}+2\alpha_{3}$ or $\beta>\alpha_{2}+\alpha_{3}+\alpha_{4}$. Since